### v1

# Zirui ${}^{\clubsuit,\spadesuit}$   Zhizhou ${}^{\vardiamond,\spadesuit}$   Zheng ${}^{\spadesuit}$   Yilin ${}^{\vardiamond,\spadesuit}$   Zhuowen ${}^{\spadesuit}$

Abstract
We present TokenCompose, a Latent Diffusion Model for text-to-image generation that achieves enhanced consistency between user-specified text prompts and model-generated images. Despite its tremendous success, the standard denoising process in the Latent Diffusion Model takes text prompts as conditions only, absent explicit constraint for the consistency between the text prompts and the image contents, leading to unsatisfactory results for composing multiple object categories. TokenCompose aims to improve multi-category instance composition by introducing the token-wise consistency terms between the image content and object segmentation maps in the finetuning stage. TokenCompose can be applied directly to the existing training pipeline of text-conditioned diffusion models without extra human labeling information. By finetuning Stable Diffusion, the model exhibits significant improvements in multi-category instance composition and enhanced photorealism for its generated images. [^1] Project done while Zirui Wang, Zhizhou Sha and Yilin Wang interned at UC San Diego. Project link: https://mlpc-ucsd.github.io/TokenCompose
Figure 1: Given a user-specified text prompt consisting of object compositions that are unlikely to appear simultaneously in a natural scene, our proposed TokenCompose method attains significant performance enhancement over the baseline Latent Diffusion Model (e.g., Stable Diffusion [56]) by being able to generate multiple categories of instances from the prompt more accurately.
1 Introduction
Despite the tremendous progress in recent text-to-image diffusion models [53, 56, 74, 58, 59] that have achieved creating images with an increasing level of quality, resolution, photorealism, and diversity, there still exists a major consistency problem between the text prompt and the generated image content. The models loose the composition capability when multiple object categories, especially those not commonly appearing simultaneously in the real world, are included in the text prompt: objects may not appear in the image or their configuration is not pleasantly good looking. Figure 1 shows examples where a state-of-the-art model, Stable Diffusion [56], fails to generate desirable image content from text prompts.
Figure 2: An overview of the TokenCompose training pipeline. Given a training prompt that faithfully describes an image, we adopt a POS tagger [1] and Grounded SAM [31, 42] to extract all binary segmentation maps of the image corresponding to noun tokens from the prompt. Then, we jointly optimize the denoising U-Net of the diffusion model with both its original denoising and our grounding objective.
Prototypical generative models of various families [30, 18, 67, 32, 55, 24, 56, 19] have reached maturity. Adding conditional training signals [50, 37, 13, 14, 53, 12, 56, 74, 27] to the generative models significantly expands their modeling capability, as well as their scope of application. In the context of Latent Diffusion Models [56], one of the most commonly applied conditions, text (e.g., captions), is injected into layers of the denoising U-Net [57] via cross-attention.
However, there exists a natural discrepancy between texts (e.g., captions) that are used to train such models and texts (e.g., prompts) that are used for generation. Whereas a caption usually describes a real image faithfully, a prompt can encapsulate image features that do not match the visual scene of any real-world images. Without fine-grained training objectives for its conditioned text, a text-to-image diffusion model often fails to generalize to arbitrary compositions that lie in the prompts [66]. This can be due to the fact that the denoising objective in text-to-image LDM is only optimized to predict the noise given a text prompt, leaving the text condition only as a facilitation in optimizing the denoising function.
By imposing training objectives that operate at the token level in the conditional text, the diffusion model learns what each token from the text means in the context of the image in an atomistic manner. Subsequently, it can be better at composing different combinations of words, phrases, etc., during inference. However, obtaining the “ground-truth” labels (e.g., segmentation maps) by humans for the corresponding text tokens is label intensive and expensive, especially for the text-image pairs [61] used to train large-scale generative models. Thanks to the recent progress in vision foundation models, Grounding DINO [42] and Segment Anything (SAM) [31], grounding segmentation maps for text tokens can be readily attained automatically.
To this end, we seek to mitigate the composition problem by developing a new algorithm, TokenCompose, which leverages models pretrained with image understanding tasks [31, 42, 1] to provide token-level training supervisions to a text-to-image generative model. We show that, by augmenting each noun token from the text prompt of a text-to-image model with segmentation grounding objectives with respect to its respective image during training, the model exhibits significant improvement in object accuracy [17], multi-category instance composition, enhanced photorealism [21] with no additional inference cost for its generated images. Along with our proposed training framework for text-to-image generative models, we also present the MultiGen benchmark, which examines the capabilities of a text-to-image generative model to compose multiple categories of instances in a single image.
2 Related Works
Compositional Generation. Efforts that aim to improve compositional image generation for text-conditioned image generative models have been focused on both the training and inference stages. One approach to improving the compositional generation of diffusion models in training is by introducing additional modules, such as a ControlNet, to specify high-level features within the image [74, 75]. However, the modules added to the diffusion models increase the size of the model, leading to additional training and inference costs. Another approach through training is to leverage a reward function to encourage faithful generation of images based on compositional prompts [25, 71, 33, 4]. Albeit their efficacy, reward functions are sparse and do not provide dense supervision signals.
Inference-based methods aim to alter the latent and/or cross-attention maps. Composable Diffusion [41] decomposes a compositional prompt into sub-prompts to generate different latents, and uses a score function to combine the latents together, while Layout Guidance Diffusion [7] uses user-defined tokens and bounding boxes to backpropagate gradients to the latent, and steer the cross-attention maps to focus on specified regions for specified tokens. Other methods apply Gaussian kernels [6] or leverage linguistic features [16, 54] to manipulate the cross-attention map. While these methods do not require further training, they add considerable cost during inference, making it cost $\times$3.37 times longer to generate a single image at most, when other generation configurations are kept the same.
Our framework is training-based and does not require additional modules to be incorporated into the image generation pipeline. Furthermore, optimizing attention maps based on segmentation maps provides dense and interpretable supervision. As a training-based method that jointly optimizes token-image correspondence and image generation, the model does not require inference-time manipulations, yet it achieves strong compositionality and competitive image quality for conditional generation.
Benchmarks for Compositional Generation. Several benchmarks have been proposed to evaluate text-conditioned image generative models for compositionality. Most methods evaluate compositionality by binding attributes to and specifying relations between objects. For example, spatial relationships examine whether two different objects appear in the correct spatial layout according to the prompt [17, 2, 25, 10, 59]. Color binding examines whether a text-conditioned image generative model can correctly assign the specified colors to different objects, especially when color assignments are counterintuitive [16, 6, 2, 25, 59]. Count binding examines whether a specified instance appears with the right number of counts specified in the prompt [59, 2, 10]. There are additional types of attribute bindings and relation specifications that are evaluated in different benchmarks, such as action and size [2], and shape and texture [25].
However, the majority of such benchmarks confound evaluating capabilities of binding the correct attributes or specifying the correct inter-object relationship with the successful generation of specified objects mentioned in the prompt, making it difficult to evaluate whether improvements are made by stronger attribute assignment & relation specification capabilities or a higher object accuracy [17, 10, 22]. VISOR [17] decouples object accuracy from spatial relationship compositionality by calculating the successful rate of correct spatial relationships conditioned on the successful generation of all specified instances from the prompt. In contrast, almost all other benchmarks do not dissociate these factors and instead evaluate compositionality on a holistic basis.
Existing benchmarks on multi-category instance composition focus on successful generation of mostly two categories. On the other hand, leading image generative models have achieved significant improvements in multi-category instance composition [3, 62], which are capable of generating multiple categories of instances with a high success rate. To fill in this research gap in evaluating multi-category instance composition beyond two categories, and to evaluate our training framework in multi-category instance composition, we propose MultiGen benchmark, which contains text prompts where each prompt contains objects from an arbitrary combination of multiple categories.
Generative Models for Image Understanding. There has been an upward trend in the use of text-conditioned image generative models for open-vocabulary image understanding tasks, such as classification [34, 11], detection [36, 8], and segmentation [38, 70, 65, 29, 45, 49, 68, 69, 64].
An inherent advantage of using generative models for image understanding tasks is being open-vocabulary, as text-to-image diffusion models are trained with open-vocabulary textual prompts. Although results using Stable Diffusion [56] for unsupervised zero-shot image understanding show great potential among methods in the same setting (i.e., zero-shot and open-vocabulary), there still exists a performance gap between these approaches and those that use specialist models for image understanding [73, 5, 42, 31, 9]. This gives us the empirical basis to optimize a generative model (e.g., Stable Diffusion) with knowledge from an understanding model. Furthermore, by optimizing a diffusion model with our approach, we also observe improved performance for downstream segmentation tasks [64, 39], which further reflects successful transfer of this knowledge.
3 Method
3.1 Preliminaries
Diffusion models [24] are widely used in conditional image generative tasks. Given an image $x\in\mathbb{R}^{H\times W\times 3}$, a normally distributed variable $\epsilon$ (e.g., noise) is added to the image with a variable extent based on a timestep $t$. Given a denoising function parameterized by a neural network $\theta$, a noisy image $x_{t}$, and a timestep $t$ uniformly sampled from $\{1,\dots,T\}$, the denoising function learns to predict the noise, $\epsilon$, following the objective (Eq. 1):
$\mathcal{L}_{DM}=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\Big{]}\,$ (1)
To improve the efficiency and control of diffusion models, two changes are performed on the original diffusion recipe – forming the Latent Diffusion Model [56].
First, instead of learning a denoising function in the image space, an image is encoded in a latent state $z_{0}=\mathcal{E}(x_{0})$ using a variational autoencoder (VAE) [30]. A random noise $\epsilon$ is added to the latent $z_{0}$, resulting in a noisy latent $z_{t}$. The training process involves computing the loss between the predicted noise $\epsilon_{\theta}$ and the ground truth noise $\epsilon$ to optimize the denoising function.
Second, a conditioning mechanism is added to the denoising function to steer the diffusion process for controllable image generation via cross-attention. In our setting, the condition $y$ is a text prompt that describes the image. To use the text via cross-attention, each token is transformed into an embedding $\tau_{\theta}(y)$ using a pretrained text encoder [52, 26]. The following shows the denoising objective for an LDM (Eq. 2):
$\mathcal{L}_{LDM}:=\mathbb{E}_{\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))\|_{2}^{2}\Big{]}\,$ (2)
We only optimize the denoising function $\epsilon_{\theta}$, which is parameterized by a U-Net [57] architecture during training. $\mathcal{E}$, $\mathcal{D}$, and $\tau_{\theta}$ are kept frozen.
3.2 Token-level Attention Loss
Consider a text prompt that is transformed into text embeddings of length $L_{\tau_{\theta}(y)}$. As $\mathcal{L}_{LDM}$ only optimizes the function so that it predicts the noise and reconstructs the image latent by removing the noise, the relationship between each token’s embedding $e_{i}$, $i\in\{1,\dots,L_{\tau_{\theta}(y)}\}$ and a noisy image latent $z_{t}$ is not optimized explicitly. This leads to a poor token-level understanding in an LDM, which can be visualized via activations of the multihead cross-attention map (i.e., $\mathcal{A}$) between the token’s embedding (i.e., $K\in\mathbb{R}^{H\times L_{\tau_{\theta}(y)}\times d_{k}}$), and the noisy image latent (i.e., $Q\in\mathbb{R}^{H\times L_{z_{t}}\times d_{k}}$). For each cross-attention layer $m\in M$ with variable latent representation resolutions $L_{z_{t}}$ in the U-Net, the cross-attention map (i.e., $\mathcal{A}\in\mathbb{R}^{L_{z_{t}}\times L_{\tau_{\theta}(y)}}$) is calculated as the following (Eq. 3 and 4):
$Q^{(h)}=W^{(h)}_{Q}\cdot\varphi(z_{t}),\;K^{(h)}=W^{(h)}_{K}\cdot\tau_{\theta}(y)\;$ (3)
$\mathcal{A}=\frac{1}{H}\sum_{h}^{H}\text{softmax}\left(\frac{Q^{(h)}(K^{(h)})^{T}}{\sqrt{d_{k}}}\right)$ (4)
where $h\in\{1,\dots,H\}$ represents each head in the multihead cross-attention, $\varphi$ is a function that flattens a two-dimensional image latent into one dimension, and $d_{k}$ is the dimension of $K$.
Figure 3: Illustration of $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$. We illustrate how $\mathcal{L}_{\text{token}}$ (bottom-right) and $\mathcal{L}_{\text{pixel}}$ (upper-right) are calculated given a cross-attention map $\mathcal{A}_{i}$ and a binary segmentation mask $\mathcal{M}_{i}$. $\mathcal{L}_{\text{token}}$ aggregates attention activations toward non-masked regions of $\mathcal{M}_{i}$, and this objective is normalized by the total activations of $\mathcal{A}_{i}$. However, it does not constrain where activations should be once inside the non-masked region. $\mathcal{L}_{\text{pixel}}$ gives precise supervision whether a pixel belongs to the segmented region, constraining where activations should be with binary values. However, it is not normalized by the total activations of $\mathcal{A}_{i}$. Combining $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$, we take advantage of the benefit of each objective while minimizing their side effects to a minimum level. We show examples of cross-attention activations from models optimized with $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$, either of them, and neither of them in Figure 4.
Empirically, we observe that training a diffusion model with only $\mathcal{L}_{LDM}$ often causes activations of cross attention maps of distinct instance tokens to fail to focus on its corresponding instance appeared in the image during training, which, in turn, results in poor capabilities in composing multiple categories of instances during inference.
To alleviate this issue for better multi-category instance composition, we add a training constraint that supervises activation regions of the cross-attention maps. Specifically, for each text token $i$ that belongs to a noun within the text caption, we acquire the binary segmentation map $\mathcal{M}_{i}$ from its respective image by leveraging foundation models trained for image understanding [42, 31]. Because cross-attention maps at different layers $m$ of the U-Net have different resolutions, we downscale the resolution of $\mathcal{M}_{i}$ to match the dimensions of its corresponding $\mathcal{A}_{i}^{(m)}$ with bilinear interpolation, followed by binarization of all values to form $\mathcal{M}_{i}^{(m)}$. Different from Layout Diffusion [7], which uses user-defined bounding boxes during inference for gradient-based guidance, we directly apply a loss function $\mathcal{L}_{\text{token}}$ that aggregates activations of cross-attention toward predicted spatial regions $\mathcal{B}_{i}=\{u\in\mathcal{M}_{i}\mid u=1\}$ jointly with $\mathcal{L}_{LDM}$ during model training. For any layer $m$, $\mathcal{L_{\text{token}}}$ is defined as follows (Eq. 5):
$\mathcal{L}_{\text{token}}=\frac{1}{N}\sum_{i}^{N}\left(1-\frac{\sum_{u\in\mathcal{B}_{i}}^{L_{z_{t}}}\mathcal{A}_{(i,u)}}{\sum_{u}^{L_{z_{t}}}\mathcal{A}_{(i,u)}}\right)^{2}$ (5)
where $\mathcal{A}_{(i,u)}\in\mathbb{R}$ represents the scalar attention activation at a spatial location $u$ of $\mathcal{A}_{i}\in\mathbb{R}^{L_{z_{t}}}$ for the cross-attention map formed by the latent and the $i$th token’s embedding. Whereas the previous approach [7] calculates the loss on each attention head of a multihead cross-attention module separately, we calculate the loss on the average of cross-attention activations on all heads (see Eq. 4). We find that the latter approach encourages different heads to activate in distinct regions of the cross-attention map, which slightly improves compositional performance and image quality. We add a scaling factor $\lambda$ to $\mathcal{L_{\text{token}}}$ with respect to $\mathcal{L}_{LDM}$ such that sufficient token-level gradients can be used to optimize token-image consistency while the denoising objective is minimally compromised.
3.3 Pixel-level Attention Loss
Although $\mathcal{L_{\text{token}}}$ substantially aggregates activations of cross-attention maps toward the target regions, a side effect of this aggregation is that the model tends to overly aggregate its activations of the cross-attention map into certain subregions of its target regions. This can be reflected by visually inspecting its cross-attention map during inference (see Figure 4) and an increased binary cross-entropy loss of cross-attention map activations $\mathcal{A}$ with respect to the target binary segmentation map $\mathcal{M}$. To overcome this problem, we use $\mathcal{L}_{\text{pixel}}$ to counteract. Formally, for a cross-attention map $\mathcal{A}$ in any layer $m$ optimized with $\mathcal{L_{\text{token}}}$, we add the pixel-level cross-entropy objective that is defined as the following (Eq. 6):
$\mathcal{L}_{\text{pixel}}$ $=-\frac{1}{L_{\tau_{\theta}(y)}L_{z_{t}}}\sum_{i}^{L_{\tau_{\theta}(y)}}\sum_{u}^{L_{z_{t}}}\left(\mathcal{M}_{(i,u)}\log(\mathcal{A}_{(i,u)}\right)$ (6)
$\quad+\left(1-\mathcal{M}_{(i,u)}\right)\log\left(1-\mathcal{A}_{(i,u)}\right)$
Figure 4: Impact on cross-attention activations with different objectives. We firstly demonstrate that finetuning the Stable Diffusion with only $\mathcal{L}_{LDM}$ does not improve grounding capabilities as much. Adding $\mathcal{L_{\text{pixel}}}$ alone causes increased cross-attention activations in general. Adding $\mathcal{L_{\text{token}}}$ plays a vital role in improving token grounding, but leads activations to aggregate in subregions of the targets. By combining $\mathcal{L_{\text{token}}}$ and $\mathcal{L_{\text{pixel}}}$, the model shows substantial improvement in grounding text tokens with image features. In this illustration, we apply the null text inversion [48] technique to all models, allowing them to generate the same image for comparable cross-attention maps. In addition, we decompose cross-attention map visualizations by different layers, attention heads, and denoising timesteps, and the results are shown in Appendix E.
We add a scaling factor $\gamma$ so that $\mathcal{L}_{\text{pixel}}$ is kept roughly constant while the model is jointly optimized by $\mathcal{L}_{LDM}$ and $\mathcal{L_{\text{token}}}$. We provide a token- and pixel-level optimization illustration in Figure 3 to demonstrate different levels of granularity of $\mathcal{L_{\text{token}}}$ and $\mathcal{L_{\text{pixel}}}$. Finally, at any single optimization step, the training objective is as follows (Eq. 7):
$\mathcal{L}_{\textsc{TokenCompose}}=\underbrace{\mathcal{L}_{LDM}}_{\text{denoise}}+\sum_{m}^{M}\left(\underbrace{\lambda\mathcal{L}_{\text{token}}^{(m)}}_{\text{token grouding}}+\underbrace{\gamma\mathcal{L}_{\text{pixel}}^{(m)}}_{\text{pixel grounding}}\right)$ (7)
4 Experiment
Multi-category Instance Composition ()normal-↑($\uparrow$)( ↑ ) Photorealism ()normal-↓($\downarrow$)( ↓ ) Eff. ()normal-↓($\downarrow$)( ↓ )
Method Object COCO Instances ADE20K Instances FID FID Latency
Accuracy MG2 MG3 MG4 MG5 MG2 MG3 MG4 MG5 (C) (F)
SD [56] 29.86 90.721.33 50.740.89 11.680.45 0.880.21 89.810.40 53.961.14 16.521.13 1.890.34 20.88 71.46 7.540.17
Composable [41] 27.83 63.330.59 21.871.01 3.250.45 0.230.18 69.610.99 29.960.84 6.890.38 0.730.22 - 75.57 13.810.15
Layout [7] 43.59 93.220.69 60.151.58 19.490.88 2.270.44 96.050.34 67.830.90 21.931.34 2.350.41 - 74.00 18.890.20
Structured [16] 29.64 90.401.06 48.641.32 10.710.92 0.680.25 89.250.72 53.051.20 15.760.86 1.740.49 21.13 71.68 7.740.17
Attn-Exct [6] 45.13 93.640.76 65.101.24 28.010.90 6.010.61 91.740.49 62.510.94 26.120.78 5.890.40 - 71.68 25.434.89
Ours 52.15 98.080.40 76.161.04 28.810.95 3.280.48 97.750.34 76.931.09 33.921.47 6.210.62 20.19 71.13 7.560.14
Table 1: Performance of our model in comparison to baselines. We evaluate the performance based on multi-category instance composition (i.e., Object Accuracy (OA) from VISOR Benchmark [17] and MG2-5 from our MultiGen Benchmark), photorealism (i.e., FID [21] from COCO and Flickr30K Entities validation splits), and inference efficiency. All comparisons are based on Stable Diffusion 1.4.
4.1 Training Details
Dataset. To study the effectiveness of token grounding objectives, we finetune the Stable Diffusion model on a subset of COCO image-caption pairs [39]. Specifically, we first select all unique images from the Visual Spatial Reasoning [40] dataset, as these images have fewer visual-linguistic ambiguities and a greater number of different categories that appear in each image. Then, we use a CLIP model [52] to select the caption with the highest semantic similarity with respect to its corresponding image. Finally, we adopt a pretrained noun parser to parse all nouns from the captions, and leverage Grounded-SAM to generate binary segmentation masks for each noun (or noun phrase) [1, 31, 42]. The final dataset consists of $\sim$ 4526 image-caption pairs and their respective binary segmentation masks. We illustrate a high-level data and training pipeline in Figure 2. We provide visualizations of dataset samples in Appendix C.
Setup. Our main experiments are performed using Stable Diffusion v1.4 [56], a popular text-to-image diffusion model for high-quality generation. We use both $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ in addition to its original denoising objective $\mathcal{L}_{LDM}$ with a constant global learning rate of 5e-6 for 24,000 steps (32,000 for Stable Diffusion v2.1) using the AdamW optimizer [44]. We trained the entire U-Net with a batch size of 1 and 4 gradient accumulation steps on a single GPU. We apply center crop to all training images and their respective segmentation maps $\mathcal{M}$. For Stable Diffusion v1.4, the cross-attention layers are located in the U-Net encoder $U_{\text{E}}$, the middle block $U_{\text{Mid}}$, and the decoder $U_{\text{D}}$. We apply $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ to all cross-attention layers in $U_{\text{Mid}}$ and $U_{\text{D}}$, but not $U_{\text{E}}$. We provide detailed setup for the entire pipeline in Appendix A.
4.2 Main Results
Baselines. We compare our finetuned Stable Diffusion v1.4 model against several baselines: (1) Composable Diffusion [41], which decomposes the prompt into different conditions and uses a score-based mechanism to denoise the image; (2) Layout Guidance Diffusion [7], which backpropagates cross-attention map gradients to the noisy latent with user-specified object tokens and bounding boxes for spatially controllable compositional generation; (3) Structured Diffusion [16], which automatically parses the prompt into a constituency tree and manipulates cross-attention key and value for compositional generation; and (4) Attend-and-Excite [6], which applies Gaussian kernels attention maps from user-specified tokens, and uses smoothed attention maps for compositional generation.
Multi-category Instance Composition. Benchmarks that examine this compositionality focus on the successful generation of multiple categories of instances in the image mentioned in the text condition. We use an existing benchmark, VISOR [17], along with our benchmark, MultiGen, to study the model’s capability in multi-category instance composition. The VISOR benchmark obtains all unique pairwise combinations of the 80 object categories from COCO [39] and converts each pair (A, B) into a text prompt with an arbitrary spatial relationship (R) following a template “$<$A$>$ $<$R$>$ $<$B$>$”, for example, “a motorcycle to the left of an elephant.” With the images generated from such prompts, VISOR uses an open-vocabulary detector [46] to detect the presence and spatial locations of each category in the pair. Object Accuracy (OA) measures the successful rate of generating instances from both categories. VISOR also provides metrics for spatial relationships. However, since our work focuses on multi-category instance composition, we only adopt the OA metric, and report relevant numbers in Table 1.
COCO Instances ($\uparrow$) ADE20K Instances ($\uparrow$)
Component Original Modified OA ($\uparrow$) MG3 MG4 MG5 MG3 MG4 MG5 FID (C) ($\downarrow$)
Ours (Stable Diffusion 1.4) 52.15 76.161.04 28.810.95 3.280.48 76.931.09 33.921.47 6.210.62 20.19
(i) finetune $\mathcal{L}_{LDM}+\lambda\mathcal{L}_{\text{token}}+\gamma\mathcal{L}_{\text{pixel}}$ frozen 29.86 50.740.89 11.680.45 0.880.21 53.961.14 16.521.13 1.890.34 20.88
(ii) $\lambda\mathcal{L}_{\text{token}}+\gamma\mathcal{L}_{\text{pixel}}$ $\mathcal{L}_{LDM}+\lambda\mathcal{L}_{\text{token}}+\gamma\mathcal{L}_{\text{pixel}}$ $\mathcal{L}_{LDM}$ 38.02 63.211.73 19.031.28 1.880.23 62.861.41 22.171.14 3.270.46 23.04
(iii) $\lambda\mathcal{L}_{\text{token}}$ $\mathcal{L}_{LDM}+\lambda\mathcal{L}_{\text{token}}+\gamma\mathcal{L}_{\text{pixel}}$ $\mathcal{L}_{LDM}+\gamma\mathcal{L}_{\text{pixel}}$ 37.46 61.951.05 18.440.98 1.810.35 65.110.99 25.340.95 4.180.53 22.32
(iv) $\gamma\mathcal{L}_{\text{pixel}}$ $\mathcal{L}_{LDM}+\lambda\mathcal{L}_{\text{token}}+\gamma\mathcal{L}_{\text{pixel}}$ $\mathcal{L}_{LDM}+\lambda\mathcal{L}_{\text{token}}$ 49.85 71.611.06 24.941.24 2.830.62 75.370.90 32.911.53 5.580.46 20.60
$U_{E}^{16\times 16}$, $U_{\text{Mid}}^{8\times 8}$, $U_{D}^{16\times 16}$ 42.92 64.421.08 18.781.03 1.470.32 67.240.98 24.771.01 3.600.57 20.45
layers w. $U_{\text{Mid}}^{8\times 8}$, $U_{D}^{16\times 16}$, $U_{D}^{16\times 16}$ 41.66 66.171.29 20.291.20 1.980.48 67.531.03 25.451.24 3.870.45 20.66
$\lambda\mathcal{L}_{\text{token}}+\gamma\mathcal{L}_{\text{pixel}}$ $U_{D}^{32\times 32}$, $U_{D}^{64\times 64}$ $U_{\text{Mid}}^{8\times 8}$, $U_{D}^{16\times 16}$ 44.27 65.471.42 19.161.07 1.590.34 69.331.33 26.801.52 4.050.68 20.75
(v) $U_{\text{Mid}}^{8\times 8}$, $U_{D}^{16\times 16}$, $U_{D}^{32\times 32}$ 49.89 73.801.33 26.281.00 2.760.35 75.491.02 33.471.27 5.870.80 20.45
Table 2: Ablation studies. We show how different objectives and cross-attention layers with $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ affect the multi-category instance compositon and photorealism. Qualitative visualizations of the cross-attention maps for (i) - (iv) are in Figure 4.
MultiGen uses a similar evaluation strategy compared to VISOR, but is designed to be a more challenging metric for multi-category instance composition. Specifically, given a set of distinct instance categories of size $N$, we randomly sample 5 categories (e.g, A, B, C, D, E), format them into a sentence (i.e., A photo of A, B, C, D, and E.), and use them as the condition for a text-to-image diffusion model to generate the image. Then, we use a strong open-vocabulary detector [47] to detect the presence of these categories in the generated image. We perform the sampling process for 80 categories of COCO instances [39] and 100 categories of ADE20K instances [76, 77] 1,000 times, resulting in 1,000 text prompts as multi-category instance combinations from each dataset. For each generated image, we leverage the detector to detect how many categories of instances appear in the image. We aggregate the overall success rate of generating 2-5 specified categories out of 5 as MG2-5.
Compositional image generation often involves inference variance. To account for this, each prompt in MultiGen is used to generate 10 rounds of images, which results in 10 $\times$ 1000 images being generated for each dataset’s category combinations. We calculate MG2 through 5 for each round and report the mean and standard deviation (in subscript) of the MG2-5 success rate out of 10 rounds in Table 1. Based on the evaluation, our model exceeds all baselines in object accuracy and MG2-4. We find that Attend-and-Excite [6] has a considerable success rate in generating all 5 categories, but it falls short of generating 2-4 categories compared to our model. We conjecture that this is due to the training distribution where captions do not include as many as 5 or more categories of instances, which inevitably leads to diminishing improvements in multi-category instance composition. In addition to aggregated results, we provide further analysis of success rate of difference categories of instances in Appendix F.
Photorealism. We compare the image quality generated from the baselines and our model using the Fréchet Inception Distance (FID) metric [21]. For baselines that do not require additional conditional input other than captions, we calculate the FID score based on 10,000 image-caption pairs sampled from the COCO validation set (C). We also report the FID metric based on 1,000 image-caption pairs from the validation set of Flickr30K entities (F) [51]. As this dataset provides labels and bounding boxes for entities in the captions [72], it can be used by inference-based methods that require such input. We report all applicable scores in Table 1. We also qualitatively evaluate image quality in conjunction with multi-category instance composition across different baselines compared to our model in Figure 5 and in Appendix D. For fairness, we use the same initial latent in each comparison.
Efficiency. As a training-based method, our model does not require additional inference-time manipulations compared to a standard generation pipeline of a text-conditioned diffusion model. On the other hand, the majority of inference-based methods impose a non-ignorable compute burden for compositional generation, where the slowest baseline, Attend-and-Excite [6], takes more than 3$\times$ of time to generate a single image. We report efficiency results in Table 1 in seconds needed to generate an image on a single NVIDIA RTX 3090 GPU with 50 DDIM steps [63] and classifier-free guidance [23].
4.3 Generalization
We evaluate whether our training method generalizes to different variants of text-to-image models. To this end, we apply $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ in addition to $\mathcal{L}_{LDM}$ to Stable Diffusion v2.1, and compare the results in multi-category instance composition and photorealism between a frozen baseline and a baseline trained only with $\mathcal{L}_{LDM}$ in Table 3. The results show that these grounding objectives also benefit Stable Diffusion v2.1 remarkably.
Multi-category Instance Composition ()normal-↑($\uparrow$)( ↑ )
Model COCO Instances ADE20K Instances FID ()($\downarrow$)( ↓ )
OA MG3 MG4 MG5 MG3 MG4 MG5 (C)
SD 2.1
 frozen 47.82 70.14 25.57 3.27 75.13 35.07 7.16 19.59
 ft. w. $\mathcal{L}_{LDM}$ 55.09 76.43 32.07 4.73 77.60 37.09 7.78 20.55
Ours 60.10 80.48 36.69 5.71 79.51 39.59 8.13 19.15
Table 3: Model generalization. We show multi-category instance composition and FID results when we apply our training approach to Stable Diffusion v2.1, which has a different (1) input and cross-attention map resolution, (2) text encoder [26], and (3) training schema (i.e., progressive distillation) [60].
4.4 Knowledge Transfer
By learning segmentation maps of each noun token via cross-attention with $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$, the model is expected to gain enhanced abilities to segment instances in the image. To verify that this knowledge is transferred from a segmentation model [31] to a diffusion model, we leverage the COCO-Gen dataset from DAAM [64]. For a fair comparison, we performed null text inversion [48] on all images from this dataset for each model. As the model reconstructs the images, we use DAAM’s algorithm and evaluation protocol to calculate the mIoU between the model’s cross-attention map and human-annotated segmentation maps. The results are shown in Table 4.
Table 4: Enhanced segmentation capabilities. We examine the grounded segmentaion knowledge transfer from Grounded-SAM [31, 42] to our finetuned diffusion model with DAAM [64].
Model mIoU ()($\uparrow$)( ↑ )
SD 1.4
 frozen 0.5371
 ft. w. $\mathcal{L}_{LDM}$ 0.5412
Ours 0.5876
4.5 Downstream Metrics
While TokenCompose is not explicitly optimized for binding attributes such as color, texture, shape to or specifying relations between objects, we show that, by improving the model’s capability in multi-category instance composition, we also observe quantitative improvements in downstream compositionality metrics. We evaluate our model using benchmarks proposed by T2I-CompBench [25], which uses various expert models [35, 52, 78] to judge the alignment between compositional prompt and the generated image. We show results of this benchmark in Table 5.
We believe that improvements in these metrics are due to the model having a higher chance of composing multiple categories of instances successfully, which serves as a prerequisite for assigning attributes to and relations between objects. We provide further analysis in Appendix B.
Attribute Binding ($\uparrow$) Object Relations (normal-$\uparrow$)
Model Color Shape Texture Spatial Non-Spat. Complex
SD 1.4
 frozen 0.3765 0.3576 0.4642 0.1161 0.3102 0.2795
 ft. w. $\mathcal{L}_{LDM}$ 0.4647 0.4598 0.5209 0.1326 0.3172 0.2912
Ours 0.5055 0.4852 0.5881 0.1815 0.3173 0.2937
Table 5: Improvements in downstream metrics. As successful composition of multiple categories of instances serves as the foundation for attribute binding and object relationship specification in compositional generation, our model shows quantitative improvement on relevant metrics, despite it is not optimized explicitly for these downstream metrics.
Figure 5: Qualitative comparison between baselines and our model. We demonstrate the effectiveness of our training framework in multi-category instance composition compared with a frozen Stable Diffusion Model [56], Composable Diffusion [41], Structured Diffusion [16], Layout Guidance Diffusion [7], and Attend-and-Excite [6]. The first three columns show composition of two categories that is deemed difficult to be generated from a pretrained Stable Diffusion model (due to rare chances of co-occurrence or significant difference in instance sizes in the real world). The last three columns show the composition of three categories where composing them requires understanding of visual representations of each text token.
5 Ablations
We ablate (1) incorporating different grounding objectives; and (2) layers applied with grounding objectives in training the denoising U-Net, and evaluate how different design strategies affect multi-category instance composition and photorealism. We show our ablation results in Table 2.
5.1 Grounding Objectives
We compare our model with a model trained only with $\mathcal{L}_{LDM}$ for the same number of optimization steps in (ii). We find that there is a moderate improvement in metrics related to multi-category instance composition comparing to (i), followed by a degeneration in photorealism (i.e., FID). We conjecture that training a model only with $\mathcal{L}_{LDM}$ in the COCO dataset brings an inherent advantage in improving composing multiple categories of instances, as the dataset often contains image-caption pairs where multiple categories of instances appear in the image and are encapsulated by the caption.
We then trained a model with only $\mathcal{L}_{LDM}$ and $\mathcal{L}_{\text{pixel}}$ (iii), and a model trained with only $\mathcal{L}_{LDM}$ and $\mathcal{L}_{\text{token}}$ (iv). We observe that $\mathcal{L}_{\text{pixel}}$ has little effect compared to $\mathcal{L}_{\text{token}}$ when used alone with $\mathcal{L}_{LDM}$. $\mathcal{L}_{\text{token}}$ plays a major role in improving multi-category instance composition and photorealism. However, empirically, we find that training the model with only $\mathcal{L}_{LDM}$ and $\mathcal{L}_{\text{token}}$ often leads to attention maps that overly activate in subregions of the instance (see Figure 4). We also observe that different training runs on this combination of losses lead to unstable inference performance in multi-category instance composition.
5.2 Layers with Grounding Objectives
Finally, we experiment with adding $\mathcal{L}_{\text{pixel}}$ and $\mathcal{L}_{\text{token}}$ at different layers of cross-attention of the denoising U-Net (v). We find that adding grounding objectives to the middle block and the decoder of the U-Net improves the overall performance of multi-category instance composition. Removing the constraint from the middle block or adding the constraint to the encoder degrades the performance of the model. Furthermore, for cross-attention layers at the decoder with variable resolutions, we find that the more layers optimized with $\mathcal{L}_{\text{pixel}}$ and $\mathcal{L}_{\text{token}}$, the better the performance in both multi-category instance composition and photorealism.
6 Limitation & Conclusion
Limitation. As one of the pioneering works exploring the potential to improve a text-conditioned generative model with image-token consistency using an understanding model, we only add supervision terms to noun tokens for the text prompt. While we show that this approach improves multi-category instance composition significantly, there are many more elements from the text prompts that one can leverage an understanding model to improve a generative model, such as adjectives, verbs, and/or determiners as fine-grained token-level training objectives.
Conclusion. We explore the possibility of leveraging foundation image understanding models to improve grounding capabilities of a text-conditioned generative model. Our training framework, TokenCompose, excels at multi-category instance composition with improved image quality. To facilitate research in this niche, we also propose MultiGen, a challenging benchmark that requires a model to generate multiple categories of instances in one image. As a fundamental challenge in compositional generation, we hope our training framework and benchmark can inspire future works that effectively leverage the synergy between understanding and generation to improve either or both directions.
Acknowledgement
This work is supported by NSF Award IIS-2127544. We thank Yifan Xu, Dou Kwark, and Kaiyi Huang for their valuable feedback.
Appendix
A Pipeline Setup
We provide detailed data generation, training and inference settings for the TokenCompose pipeline setup in Table 6. Unless otherwise specified, experiments are run based on this set of settings. For settings that are not reported in Table 6, we follow the default values provided by their respective codebases.
Setting Value
Data Generation
CLIP Model ViT-L/14@336px [52, 15]
POS Tagger flair/pos-english [1]
Target POS Tags NN / NNS / NNP / NNPS
G. DINO Model groundingdino_swint_ogc [42, 5, 43]
G. DINO Box Thres. 0.25
G. DINO Text Thres. 0.25
SAM Model sam_hq_vit_h [15, 31, 28]
Training
Resolution SD 1.4 $512\times 512$
Resolution SD 2.1 $768\times 768$
Image Processing Center Crop + Resize
Batch Size 1
Grad. Accum. Steps 4
Grad. Ckpting. True
Train Steps SD 1.4 24000
Train Steps SD 2.1 32000
Learning Rate 5e-6
LR Scheduler Constant
LR Warmup False
$\lambda$ for $\mathcal{L}_{token}$ 1e-3
$\gamma$ for $\mathcal{L}_{pixel}$ 5e-5
$\mathcal{L}_{token}$ and $\mathcal{L}_{pixel}$ all layers in $U_{\text{Mid}}$ and $U_{\text{D}}$
Clf-Free Guidance [23] False
Inference
Resolution SD 1.4 $512\times 512$
Resolution SD 2.1 $768\times 768$
Timestep Scheduler PNDMScheduler
# Inference Steps 50
Clf-Free Guidance [23] True
Guidance Scale 7.5
Table 6: Model choices and settings for data generation, training, and inference. We provide a comprehensive list of pipeline details for data generation (e.g., caption selection, noun extraction and segmentation map generation), training (e.g., finetuning the Stable Diffusion model), and inference (e.g., evaluation of our finetuned Stable Diffusion model).
B Conditional Downstream Metrics
As multi-category instance composition serves as a prerequisite for successful downstream text-conditioned compositional generation, we conjecture that the improvements in downstream metrics are improved by higher chances of generating all instances mentioned in the prompt. In this section, we provide an additional object accuracy metric in Table 7, and show that TokenCompose improves object accuracy in all downstream metrics being evaluated. Furthermore, in Figure 6, we demonstrate how better multi-category instance composition capabilities can improve these metrics by plotting the improvement curve from (1) a frozen Stable Diffusion to (2) a Stable Diffusion finetuned with only $\mathcal{L}_{LDM}$ objective, and finally to (3) a Stable Diffusion model finetuned with both $\mathcal{L}_{LDM}$ and our grounding objectives conditioned and unconditioned on successful generation of all instances from the compositional prompts.
From the results, we observe that the improvements in all attribute binding metrics (e.g., color, shape, texture) and the majority of object relation metrics (e.g., spatial, complex) are much more significant in the unconditional case than in the conditional case. The only downstream benchmark where the difference in improvement is insignificant is the non-spatial compositionality. We believe that this insignificance can be explained by two factors: (1) lowest amount of improvement in object accuracy for this specific downstream benchmark, as shown in Table 7 and (2) relatively low correlation between automatic scores (i.e., CLIP Score [52, 20]) and human ratings among all compositional benchmarks from the T2I-CompBench [25]; this indicates that the evaluation model may have a comparably weak discriminative capability for this specific task.
Model Color Shape Texture Spatial Non-Spat. Complex
SD 1.4
 frozen 47.40 25.33 15.27 19.40 45.10 26.60
 ft. w. $\mathcal{L}_{LDM}$ 55.70 28.49 18.95 27.05 47.33 29.33
Ours 62.92 32.95 25.59 33.30 48.94 32.10
Table 7: Object accuracy in downstream compositional metrics. We calculate the object accuracy metric (i.e., success rate of generating all instances in the prompt based on a detection model [47]) for each of the compositional benchmarks from T2I-CompBench [25].
Figure 6: Improvement comparison of unconditional and conditional downstream compositional metrics. We illustrate that a significant margin of downstream compositional metrics is improved due to enhanced capabilities of multi-category instance composition. In this figure, we calculate the evaluation metrics from T2I-CompBench [25] by conditioning on the successful generation of all instances mentioned in the prompts, and compare the amount of quantitative improvement with the same metrics that are not conditioned on successfull generation of all instances.
C Grounded COCO Dataset
As seen in Figure 2, we adopt a POS Tagger and the Grounded SAM [1, 42, 31] to extract binary segmentation maps from noun tokens from the image-text pair dataset. In this section, we aim to expand the visualizations of the generated data in Figure 8.
Each row and column in Figure 8 represents a single data that the model is trained with. From the left to the right of the data are the caption, the input image, and the grounded binary segmentation maps.
The bold and underlined text in the caption represents noun tokens captured by the POS tagger where their corresponding segmentation maps are extracted from the Grounded SAM. Italicized and red text in the caption represents noun tokens captured by the POS tagger but does not have segmentation maps extracted from the Grounded SAM due to the model not being able to find the corresponding objects.
Grounded segmentation maps on the right are paired with their correpsonding tokens. For tokens with a green background, their segmentation maps successfully capture the aligned contents in the image, whereas for a small fraction of tokens with an orange background, their segmentation maps do not capture the aligned contents in the image.
We also provide reference on approximations (i.e., word split by space) of (1) average number of tokens per caption, (2) average number of noun tokens per caption, and (3) average number of noun tokens that have their corresponding segmentation maps per caption in Figure 7. As shown in the figure, our training dataset contains an average of 3.71 noun tokens overall and 3.21 noun tokens that have their corresponding segmentation maps.
Figure 7: Approximations of the average number of tokens with grounding objectives per training prompt. From left to right, we show the distribution of number of tokens per caption, number of noun tokens per caption, and number of noun tokens that have generated segmentation maps per caption.
Figure 8: Grounding dataset visualization. We provide visualizations of 50 selected image-text pairs (number of noun tokens with segmentation maps $\geq 5$) and their corresponding token-level binary segmentation maps from the training data to facilitate understanding of our training dataset.
D Sampled Images from MultiGen
MultiGen benchmark plays a crucial role in evaluating multi-category instance composition capabilities of text-to-image models. Our model exhibits significant improvements in multi-category instance composition compared to previous methods, as evidenced by the quantitative results from Table 1.
To offer a more intuitive sense in this compositional task, we visualize the generated images from various prompts in this benchmark (e.g., MG COCO Instances and MG ADE20K Instances). We use the same latent for each comparison for fairness. The images are presented in Figure 9.
Figure 9: Sampled images with prompts from our MultiGen benchmark. To facilitate understanding of our multi-category instance composition benchmark as well as further qualitative comparisons among different baselines, we provide sampled images from MultiGen with COCO instances as well as ADE20K instances.
E Attention Map Visualizations
Figure 10: Cross-attention map visualizations by different cross-attention layers at denoising U-Net. We visualize the cross-attention map between a frozen Stable Diffusion [56] model and our model at the middle block and decoder layers of the denoising U-Net, where cross-attention at these layers are trained with our grounding objectives. In the upper example, we leverage null text inversion [48] to let two models reconstruct similar images using different latents for comparable cross-attention maps to demonstrate stronger grounding capabilities of our model. In the lower example, we use the same initial latent for two different models to generate images to demonstrate how stronger grounding capabilities lead to better compositionality.
Figure 11: Cross-attention map visualizations by different cross-attention heads at the denoising U-Net. We visualize the cross-attention map of each head at the last layer (i.e., $U_{\text{D}}^{64\times 64}$) of the U-Net decoder between a frozen Stable Diffusion [56] model and our model. We use the same setting as in Figure 10 but with a different prompt for a more diverse visualization. We show that our grounding objective allows flexibility of different heads to attend to different regions of the latent.
Figure 12: Cross-attention map visualizations by different denoising time steps at the denoising U-Net. We visualize the cross-attention map of the last layer (i.e., $U_{\text{D}}^{64\times 64}$) of the U-Net decoder between a frozen Stable Diffusion [56] model and our model at time step 1 and every 5 steps based on a 50-step DDIM [63] scheduler. We use the same setting as in Figure 10 with a different prompt for a more diverse visualization. We show that our grounding objective enables cross-attention of different object tokens to aggregate at different regions of the noisy latent early during inference. This enables the model to generate different categories of instances more successfully, leading to better multi-category instance composition capabilities.
To gain a better understanding of how incorporating grounding objectives into text-to-image models during training affects cross-attention maps for image reconstruction [48] & generation tasks at inference time, we provide token-level cross-attention map visualizations on three different axes: (1) different cross-attention layers with various resolutions (Figure 10); (2) different heads of the multi-head cross-attention (Figure 11; and (3) different timesteps during the denoising process (Figure 12).
F Instance Generation Success Rate
Multi-category instance composition benchmarks provide the success rate of generating all categories of instances from the evaluation prompts. To gain a deep understanding of this capability, we provide the success rate from the benchmarks that we evaluate our model with by category.
Given a set of compositional prompts, we calculate the count of each category that appears in these prompts along with images where the instance(s) of this category is/are detected by a detection model [47]. We divide the number of images that contain a specific category of the instance by the number of prompts that contains this category to acquire the success rate and report the numbers in Figure 13.
Figure 13: Instance generation success rate in multi-category instance composition benchmarks. We provide the success rate (from a 0-1 scale) of generating instances with our best-performing model for VISOR [17], MG-COCO, and MG-ADE20K benchmarks.

### v2

@@
-Zirui ${}^{\clubsuit,\spadesuit}$   Zhizhou ${}^{\vardiamond,\spadesuit}$   Zheng ${}^{\spadesuit}$   Yilin ${}^{\vardiamond,\spadesuit}$   Zhuowen ${}^{\spadesuit}$
+https://mlpc-ucsd.github.io/TokenCompose
 Abstract
@@
 Despite its tremendous success, the standard denoising process in the Latent Diffusion Model takes text prompts as conditions only, absent explicit constraint for the consistency between the text prompts and the image contents, leading to unsatisfactory results for composing multiple object categories.
-TokenCompose aims to improve multi-category instance composition by introducing the token-wise consistency terms between the image content and object segmentation maps in the finetuning stage.
+Our proposed TokenCompose aims to improve multi-category instance composition by introducing the token-wise consistency terms between the image content and object segmentation maps in the finetuning stage.
 TokenCompose can be applied directly to the existing training pipeline of text-conditioned diffusion models without extra human labeling information.
-By finetuning Stable Diffusion, the model exhibits significant improvements in multi-category instance composition and enhanced photorealism for its generated images.
-[^1] Project done while Zirui Wang, Zhizhou Sha and Yilin Wang interned at UC San Diego.
-Project link: https://mlpc-ucsd.github.io/TokenCompose
+By finetuning Stable Diffusion with our approach, the model exhibits significant improvements in multi-category instance composition and enhanced photorealism for its generated images.
+[^1] Project done while ZW, ZS and YW interned at UC, San Diego.
+Correspondence to ZW at zw1300@cs.princeton.edu
 Figure 1: Given a user-specified text prompt consisting of object compositions that are unlikely to appear simultaneously in a natural scene, our proposed TokenCompose method attains significant performance enhancement over the baseline Latent Diffusion Model (e.g., Stable Diffusion [59]) by being able to generate multiple categories of instances from the prompt more accurately.
 1 Introduction
-Despite the tremendous progress in recent text-to-image diffusion models [53, 56, 74, 58, 59] that have achieved creating images with an increasing level of quality, resolution, photorealism, and diversity, there still exists a major consistency problem between the text prompt and the generated image content.
+Despite the tremendous progress in recent text-to-image diffusion models [56, 59, 79, 61, 62, 18, 76, 7, 22] that have achieved creating images with an increasing level of quality, resolution, photorealism, and diversity, there still exists a major consistency problem between the text prompt and the generated image content.
 The models loose the composition capability when multiple object categories, especially those not commonly appearing simultaneously in the real world, are included in the text prompt: objects may not appear in the image or their configuration is not pleasantly good looking.
@@
 Generative Models for Image Understanding.
-There has been an upward trend in the use of text-conditioned image generative models for open-vocabulary image understanding tasks, such as classification [34, 11], detection [36, 8], and segmentation [38, 70, 65, 29, 45, 49, 68, 69, 64].
+There has been an upward trend in the use of text-conditioned image generative models for open-vocabulary image understanding tasks, such as classification [37, 12], detection [39, 9], and segmentation [41, 74, 68, 32, 48, 52, 71, 73, 67, 72, 81].
 An inherent advantage of using generative models for image understanding tasks is being open-vocabulary, as text-to-image diffusion models are trained with open-vocabulary textual prompts.
@@
 1):
-$\mathcal{L}_{DM}=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\Big{]}\,$ (1)
+$\mathcal{L}_{DM}=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\Big{]}\,\vspace{-1mm}$ (1)
 To improve the efficiency and control of diffusion models, two changes are performed on the original diffusion recipe – forming the Latent Diffusion Model [59].
@@
 2):
-$\mathcal{L}_{LDM}:=\mathbb{E}_{\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))\|_{2}^{2}\Big{]}\,$ (2)
+$\mathcal{L}_{LDM}:=\mathbb{E}_{\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))\|_{2}^{2}\Big{]}\,\vspace{-2mm}$ (2)
 We only optimize the denoising function $\epsilon_{\theta}$, which is parameterized by a U-Net [60] architecture during training. $\mathcal{E}$, $\mathcal{D}$, and $\tau_{\theta}$ are kept frozen.
@@
 3 and 4):
-$Q^{(h)}=W^{(h)}_{Q}\cdot\varphi(z_{t}),\;K^{(h)}=W^{(h)}_{K}\cdot\tau_{\theta}(y)\;$ (3)
-$\mathcal{A}=\frac{1}{H}\sum_{h}^{H}\text{softmax}\left(\frac{Q^{(h)}(K^{(h)})^{T}}{\sqrt{d_{k}}}\right)$ (4)
+$Q^{(h)}=W^{(h)}_{Q}\cdot\varphi(z_{t}),\;K^{(h)}=W^{(h)}_{K}\cdot\tau_{\theta}(y)\;\vspace{-2mm}$ (3)
+$\mathcal{A}=\frac{1}{H}\sum_{h}^{H}\text{softmax}\left(\frac{Q^{(h)}(K^{(h)})^{T}}{\sqrt{d_{k}}}\right)\vspace{-2mm}$ (4)
 where $h\in\{1,\dots,H\}$ represents each head in the multihead cross-attention, $\varphi$ is a function that flattens a two-dimensional image latent into one dimension, and $d_{k}$ is the dimension of $K$.
 Figure 3: Illustration of $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$.
-We illustrate how $\mathcal{L}_{\text{token}}$ (bottom-right) and $\mathcal{L}_{\text{pixel}}$ (upper-right) are calculated given a cross-attention map $\mathcal{A}_{i}$ and a binary segmentation mask $\mathcal{M}_{i}$.
+We illustrate how $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ are calculated given a cross-attention map $\mathcal{A}_{i}$ and a binary segmentation mask $\mathcal{M}_{i}$.
 $\mathcal{L}_{\text{token}}$ aggregates attention activations toward non-masked regions of $\mathcal{M}_{i}$, and this objective is normalized by the total activations of $\mathcal{A}_{i}$.
@@
 In this illustration, we apply the null text inversion [51] technique to all models, allowing them to generate the same image for comparable cross-attention maps.
-In addition, we decompose cross-attention map visualizations by different layers, attention heads, and denoising timesteps, and the results are shown in Appendix E.
 We add a scaling factor $\gamma$ so that $\mathcal{L}_{\text{pixel}}$ is kept roughly constant while the model is jointly optimized by $\mathcal{L}_{LDM}$ and $\mathcal{L_{\text{token}}}$.
@@
 4 Experiment
-Multi-category Instance Composition ()normal-↑($\uparrow$)( ↑ ) Photorealism ()normal-↓($\downarrow$)( ↓ ) Eff. ()normal-↓($\downarrow$)( ↓ )
+Multi-category Instance Composition ()($\uparrow$)( ↑ ) Photorealism ()($\downarrow$)( ↓ ) Eff. ()($\downarrow$)( ↓ )
 Method Object COCO Instances ADE20K Instances FID FID Latency
@@
 We illustrate a high-level data and training pipeline in Figure 2.
-We provide visualizations of dataset samples in Appendix C.
 Setup.
@@
 For Stable Diffusion v1.4, the cross-attention layers are located in the U-Net encoder $U_{\text{E}}$, the middle block $U_{\text{Mid}}$, and the decoder $U_{\text{D}}$.
-We apply $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ to all cross-attention layers in $U_{\text{Mid}}$ and $U_{\text{D}}$, but not $U_{\text{E}}$.
-We provide detailed setup for the entire pipeline in Appendix A.
+We apply $\mathcal{L}_{\text{token}}$ and $\mathcal{L}_{\text{pixel}}$ to all cross-attention layers in $U_{\text{Mid}}$ and $U_{\text{D}}$.
 4.2 Main Results
@@
 We use an existing benchmark, VISOR [19], along with our benchmark, MultiGen, to study the model’s capability in multi-category instance composition.
-The VISOR benchmark obtains all unique pairwise combinations of the 80 object categories from COCO [39] and converts each pair (A, B) into a text prompt with an arbitrary spatial relationship (R) following a template “$<$A$>$ $<$R$>$ $<$B$>$”, for example, “a motorcycle to the left of an elephant.” With the images generated from such prompts, VISOR uses an open-vocabulary detector [46] to detect the presence and spatial locations of each category in the pair.
+The VISOR benchmark obtains all unique pairwise combinations of the 80 object categories from COCO [42] and converts each pair (A, B) into a text prompt with an arbitrary spatial relationship (R) following a template “$<$A$>$ $<$R$>$ $<$B$>$", for example, “a motorcycle to the left of an elephant." With the images generated from such prompts, VISOR uses an open-vocabulary detector [49] to detect the presence and spatial locations of each category in the pair.
 Object Accuracy (OA) measures the successful rate of generating instances from both categories.
@@
 To account for this, each prompt in MultiGen is used to generate 10 rounds of images, which results in 10 $\times$ 1000 images being generated for each dataset’s category combinations.
-We calculate MG2 through 5 for each round and report the mean and standard deviation (in subscript) of the MG2-5 success rate out of 10 rounds in Table 1.
+We calculate MG2-5 for each round and report the mean and standard deviation (in subscript) of the MG2-5 success rate out of 10 rounds in Table 1.
 Based on the evaluation, our model exceeds all baselines in object accuracy and MG2-4.
-We find that Attend-and-Excite [6] has a considerable success rate in generating all 5 categories, but it falls short of generating 2-4 categories compared to our model.
+We find that Attend-and-Excite [6] has a considerable success rate in generating all 5 categories, but it falls short of generating 2-4 categories comparably.
 We conjecture that this is due to the training distribution where captions do not include as many as 5 or more categories of instances, which inevitably leads to diminishing improvements in multi-category instance composition.
-In addition to aggregated results, we provide further analysis of success rate of difference categories of instances in Appendix F.
 Photorealism.
@@
 We report all applicable scores in Table 1.
-We also qualitatively evaluate image quality in conjunction with multi-category instance composition across different baselines compared to our model in Figure 5 and in Appendix D.
+We also qualitatively evaluate image quality in conjunction with multi-category instance composition across different baselines compared to our model in Figure 5.
 For fairness, we use the same initial latent in each comparison.
 Efficiency.
-As a training-based method, our model does not require additional inference-time manipulations compared to a standard generation pipeline of a text-conditioned diffusion model.
+As a training-based method, our model does not require additional inference-time manipulations compared to a standard text-to-image diffusion pipeline.
 On the other hand, the majority of inference-based methods impose a non-ignorable compute burden for compositional generation, where the slowest baseline, Attend-and-Excite [6], takes more than 3$\times$ of time to generate a single image.
@@
 The results show that these grounding objectives also benefit Stable Diffusion v2.1 remarkably.
-Multi-category Instance Composition ()normal-↑($\uparrow$)( ↑ )
+Multi-category Instance Composition ()($\uparrow$)( ↑ )
 Model COCO Instances ADE20K Instances FID ()($\downarrow$)( ↓ )
@@
 We show results of this benchmark in Table 5.
-We believe that improvements in these metrics are due to the model having a higher chance of composing multiple categories of instances successfully, which serves as a prerequisite for assigning attributes to and relations between objects.
-We provide further analysis in Appendix B.
-Attribute Binding ($\uparrow$) Object Relations (normal-$\uparrow$)
+We believe that improvements in these metrics are due to the model having higher chances of composing multiple categories of instances, which serves as a prerequisite for assigning attributes to and relations between objects.
+Attribute Binding ($\uparrow$) Object Relations ($\uparrow$)
 Model Color Shape Texture Spatial Non-Spat. Complex
@@
 We find that adding grounding objectives to the middle block and the decoder of the U-Net improves the overall performance of multi-category instance composition.
-Removing the constraint from the middle block or adding the constraint to the encoder degrades the performance of the model.
+Removing the constraint from the middle block or adding the constraint to the encoder degrades the performance.
 Furthermore, for cross-attention layers at the decoder with variable resolutions, we find that the more layers optimized with $\mathcal{L}_{\text{pixel}}$ and $\mathcal{L}_{\text{token}}$, the better the performance in both multi-category instance composition and photorealism.
@@
 Acknowledgement
-This work is supported by NSF Award IIS-2127544. We thank Yifan Xu, Dou Kwark, and Kaiyi Huang for their valuable feedback.
+This work is supported by NSF Award IIS-2127544.
+We thank Yifan Xu and Dou Kwark from UC San Diego, Kaiyi Huang from University of Hong Kong, and Adithya Bhaskar and Ofir Press from the Princeton NLP Group for discussions and/or feedback.
 Appendix
-A Pipeline Setup
+7 Pipeline Setup
 We provide detailed data generation, training and inference settings for the TokenCompose pipeline setup in Table 6.
@@
 We provide a comprehensive list of pipeline details for data generation (e.g., caption selection, noun extraction and segmentation map generation), training (e.g., finetuning the Stable Diffusion model), and inference (e.g., evaluation of our finetuned Stable Diffusion model).
-B Conditional Downstream Metrics
+8 Conditional Downstream Metrics
+Figure 6: Improvement comparison of unconditional and conditional downstream compositional metrics.
+We illustrate that a significant margin of downstream compositional metrics is improved due to enhanced capabilities of multi-category instance composition.
+In this figure, we calculate the evaluation metrics from T2I-CompBench [28] by conditioning on the successful generation of all instances mentioned in the prompts, and compare the amount of quantitative improvement with the same metrics that are not conditioned on successfull generation of all instances.
 As multi-category instance composition serves as a prerequisite for successful downstream text-conditioned compositional generation, we conjecture that the improvements in downstream metrics are improved by higher chances of generating all instances mentioned in the prompt.
@@
 We calculate the object accuracy metric (i.e., success rate of generating all instances in the prompt based on a detection model [50]) for each of the compositional benchmarks from T2I-CompBench [28].
-Figure 6: Improvement comparison of unconditional and conditional downstream compositional metrics.
-We illustrate that a significant margin of downstream compositional metrics is improved due to enhanced capabilities of multi-category instance composition.
-In this figure, we calculate the evaluation metrics from T2I-CompBench [25] by conditioning on the successful generation of all instances mentioned in the prompts, and compare the amount of quantitative improvement with the same metrics that are not conditioned on successfull generation of all instances.
-C Grounded COCO Dataset
+Figure 7: Approximations of the average number of tokens with grounding objectives per training prompt.
+From left to right, we show the distribution of number of tokens per caption, number of noun tokens per caption, and number of noun tokens that have generated segmentation maps per caption.
+Figure 8: More samples in multi-category instance composition.
+Figure 9: Downstream applications in prompt-to-prompt [22] image editing and zero-shot outpainting and inpainting.
+Figure 10: Categories with poor & strong performance
+Figure 11: Failure case analysis in multi-category compositionality.
+Figure 12: Sampled images with prompts from our MultiGen benchmark.
+To facilitate understanding of our multi-category instance composition benchmark as well as further qualitative comparisons among different baselines, we provide sampled images from MultiGen with COCO instances as well as ADE20K instances.
+Figure 13: Grounding dataset visualization.
+We provide visualizations of 50 selected image-text pairs (number of noun tokens with segmentation maps $\geq 5$) and their corresponding token-level binary segmentation maps from the training data to facilitate understanding of our training dataset.
+9 More Examples
+Multi-category Instance Composition.
+We provide more visualizations in Figure.
+8 to illustrate capabilities of TokenCompose in multi-category instance composition.
+In addition to being able to generate multiple categories of instances successfully, we show that object affordance (e.g., attach, sit, support, etc) can be maintained, which indicates that TokenCompose is able to implicitly learn the basic “physical rules" via the object token and segmentation consistency constraints.
+Downstream Applications. We show in Figure. 9 positive impacts in downstream applications. As expected, TokenCompose shows its effectiveness in tasks beyond text-to-image generation.
+Benchmark.
+MultiGen benchmark plays a crucial role in evaluating multi-category instance composition capabilities of text-to-image models.
+To offer a more intuitive sense in this compositional task, we visualize the generated images from various prompts in this benchmark (e.g., MG COCO Instances and MG ADE20K Instances).
+We use the same latent for each comparison for fairness.
+The images are presented in Figure 12.
+10 Grounded COCO Dataset
 As seen in Figure 2, we adopt a POS Tagger and the Grounded SAM [1, 45, 34] to extract binary segmentation maps from noun tokens from the image-text pair dataset.
-In this section, we aim to expand the visualizations of the generated data in Figure 8.
-Each row and column in Figure 8 represents a single data that the model is trained with.
+We aim to expand the visualizations of the generated data in Figure 13.
+Each row and column represents a single data that the model is trained with.
 From the left to the right of the data are the caption, the input image, and the grounded binary segmentation maps.
 The bold and underlined text in the caption represents noun tokens captured by the POS tagger where their corresponding segmentation maps are extracted from the Grounded SAM.
-Italicized and red text in the caption represents noun tokens captured by the POS tagger but does not have segmentation maps extracted from the Grounded SAM due to the model not being able to find the corresponding objects.
+Italicized and red text in the caption represents noun tokens captured by the POS tagger but does not have segmentation maps extracted from the Grounded SAM due to the model not being able to locate the objects.
 Grounded segmentation maps on the right are paired with their correpsonding tokens.
@@
 As shown in the figure, our training dataset contains an average of 3.71 noun tokens overall and 3.21 noun tokens that have their corresponding segmentation maps.
-Figure 7: Approximations of the average number of tokens with grounding objectives per training prompt.
-From left to right, we show the distribution of number of tokens per caption, number of noun tokens per caption, and number of noun tokens that have generated segmentation maps per caption.
-Figure 8: Grounding dataset visualization.
-We provide visualizations of 50 selected image-text pairs (number of noun tokens with segmentation maps $\geq 5$) and their corresponding token-level binary segmentation maps from the training data to facilitate understanding of our training dataset.
-D Sampled Images from MultiGen
-MultiGen benchmark plays a crucial role in evaluating multi-category instance composition capabilities of text-to-image models.
-Our model exhibits significant improvements in multi-category instance composition compared to previous methods, as evidenced by the quantitative results from Table 1.
-To offer a more intuitive sense in this compositional task, we visualize the generated images from various prompts in this benchmark (e.g., MG COCO Instances and MG ADE20K Instances).
-We use the same latent for each comparison for fairness.
-The images are presented in Figure 9.
-Figure 9: Sampled images with prompts from our MultiGen benchmark.
-To facilitate understanding of our multi-category instance composition benchmark as well as further qualitative comparisons among different baselines, we provide sampled images from MultiGen with COCO instances as well as ADE20K instances.
-E Attention Map Visualizations
+11 Analysis
+Attention Visualizations.
+To gain a better understanding of how incorporating grounding objectives into text-to-image models during training affects cross-attention maps for image reconstruction [51] & generation tasks at inference time, we provide token-level cross-attention map visualizations on three axes: (1) different cross-attention layers with various resolutions (Figure 14); (2) different heads of the multi-head cross-attention (Figure 15; and (3) different timesteps during the denoising process (Figure 16).
+Multi-category Instance Composition Success Rate.
+Given a set of compositional prompts, we calculate the count of each category that appears in these prompts along with images where the instance(s) of this category is/are detected by a detection model [50].
+We divide the number of images that contain a specific category of the instance by the number of prompts that contains this category to acquire the success rate and report the numbers in Figure 17.
+Failure Cases.
+We provide generated and detected samples for categories with low and high success rates in Figure.
+10.
+We believe that one explanation for the poor performance categories have variants and viewpoints with drastic visual differences, so learning and generating them is harder.
+Further, we aggregate patterns of common failures in our multi-category instance composition in Figure.
+11.
+We believe that visual commonsense reasoning aspect of the generative model would be an area of improvement.
 Figure 14: Cross-attention map visualizations by different cross-attention layers at denoising U-Net.
@@
 This enables the model to generate different categories of instances more successfully, leading to better multi-category instance composition capabilities.
-To gain a better understanding of how incorporating grounding objectives into text-to-image models during training affects cross-attention maps for image reconstruction [48] & generation tasks at inference time, we provide token-level cross-attention map visualizations on three different axes: (1) different cross-attention layers with various resolutions (Figure 10); (2) different heads of the multi-head cross-attention (Figure 11; and (3) different timesteps during the denoising process (Figure 12).
-F Instance Generation Success Rate
-Multi-category instance composition benchmarks provide the success rate of generating all categories of instances from the evaluation prompts.
-To gain a deep understanding of this capability, we provide the success rate from the benchmarks that we evaluate our model with by category.
-Given a set of compositional prompts, we calculate the count of each category that appears in these prompts along with images where the instance(s) of this category is/are detected by a detection model [47].
-We divide the number of images that contain a specific category of the instance by the number of prompts that contains this category to acquire the success rate and report the numbers in Figure 13.
 Figure 17: Instance generation success rate in multi-category instance composition benchmarks.